OpenStudy (anonymous):

Compute f '(a) algebraically for the given value of a f(x) = x2 − 9; a = 3

OpenStudy (anonymous):

If f(x) = x^2 - 9, then f'(x) = 2x So at a, f'(3) = 2(3) = 6

OpenStudy (anonymous):

Awesome thank you if I had to compute the derivative this way f(a+h)-f(a)/h what would the work look like?

OpenStudy (anonymous):

Yes, but go the easy way unless you are specifically asked to use the definition of the derivative (what you suggested).

OpenStudy (anonymous):

Absolutely I agree, however lets say my teacher wants me to explain in this form: f(a+h)-f(a)/h using the equation confuses me but I need to know it unfortunately. Do you know how to do the problem that way? if you could I appreciate it. Thank you